# ABACUS: An Electronic Structure Analysis Package for the AI Era

### v1

@@ IV.3.1 Formulas
The stochastic density functional theory (sDFT) Baer, Neuhauser, and Rabani (2013) was proposed to circumvent the $O(N^{3})$ scaling of KS-DFT calculations. In this method, by introducing stochastic orbitals and the Chebyshev expansion scheme, the electron density can be evaluated by tracing operators, bypassing the diagonalization of the Hamiltonian. Since the computational costs of tracing operations scale linearly with system size, this approach significantly enhances computational efficiency for large systems. Later, the finite-temperature sDFT Cytter et al. (2018) defines the Fermi-Dirac operator at finite temperature as
@@
However, the use of stochastic orbitals inevitably introduces stochastic errors. To achieve higher accuracy, a large number of stochastic orbitals are typically required, which substantially increases the computational time despite its linear scaling property. For instance, fragmentation approaches have been developed to compute molecular systems Neuhauser, Baer, and Rabani (2014) and covalent materials Arnon et al. (2017); Chen et al. (2019a), while stochastic “embedding” methods have been used to calculate p-nitroaniline in water Li et al. (2019). Additionally, the energy window method proposed by Chen et al. Chen et al. (2019b) decomposes the electron density into different components based on orbital energies to reduce stochastic errors, demonstrating good generality and robustness. For simulations of materials at extremely high temperatures, mixed stochastic-deterministic Density Functional Theory (mDFT) was proposed White and Collins (2020). The method combines the advantages of traditional KS-DFT and sDFT. The mDFT method retains the computational efficiency of sDFT for high-temperature systems while reducing stochastic errors using some of the deterministic Kohn-Sham orbitals.
@@
$=2\mathrm{Tr}\left[\hat{f}_{H}|\mathbf{r}\rangle\langle\mathbf{r}|\right]$
$=2\mathrm{Tr}\left[\hat{f}_{H}^{1/2}|\mathbf{r}\rangle\langle\mathbf{r}|\hat{f}_{H}^{1/2}\right],$
where the chemical potential $\mu$ in $\hat{f}_{H}$ is determined by solving the conservation of electron number via
@@
IV.3.2 Stochastic Orbitals
@@
$\langle\phi_{j}|\chi_{a}\rangle=\frac{1}{\sqrt{N\chi}}\exp(i2\pi\theta_{j}^{a}),$ (52)
@@
|$\ket{\tilde{\chi}_{a}}=\ket{\chi_{a}}-\sum_{i=1}^{N_{\varphi}}\langle\varphi_{i}|\chi_{a}\rangle\ket{\varphi_{i}},$ (56)
@@
For example, in mDFT, the formula for electron density is
@@
Note that here we use a finite number of stochastic orbitals $N_{\chi}$, which introduces stochastic errors.
IV.3.3 Plane-Wave-Based Implementation
We have implemented sDFT and mDFT methods based on plane wave basis Liu and Chen (2022b) and periodic boundary conditions in ABACUS. Furthermore, both methods can be use with the $k$-point sampling method. The flowchart of sDFT and mDFT is shown in Fig. 14.
Figure 14: Flowchart of mDFT and sDFT in ABACUS. When the number of KS orbitals $N_{\phi}$ is larger than 0, it is the flowchart of mDFT. When the number of KS orbitals $N_{\phi}$ reduces to 0, indicating that the dashed box does not exist, it degenerates into the flowchart of sDFT.
@@ IV.3.3 Plane-Wave-Based Implementation
$\rho(\mathbf{r})=2\sum_{\mathbf{k}}W(\mathbf{k})\sum_{a,\mathbf{G}}\left|\left\langle\mathbf{r}|\mathbf{k+G}\right\rangle\left\langle\mathbf{k+G}\left|\hat{f}_{H}^{1/2}\right|\chi_{a\mathbf{k}}\right\rangle\right|^{2}.$ (61)
If we define
@@
the electron density can be written as
@@
For the force, only the nonlocal potential part is different from the traditional Kohn-Sham method, which is given by
$\mathbf{F}_{I,\tau}^{\mathrm{NL}}=$ $2{\rm Tr}\Big{[}-i\hat{f}_{H}\sum_{\mathbf{GG^{\prime}}}(\mathbf{G^{\prime}-G})e^{i(\mathbf{G^{\prime}-G})\cdot\mathbf{R}_{I}}$ (64)
@@
$=$ $-2i\sum_{\mathbf{k}}W(\mathbf{k})\sum_{a{\mathbf{G,G^{\prime}}}}\gamma_{a\mathbf{k}}(\mathbf{G})\gamma_{a\mathbf{k}}(\mathbf{G^{\prime}})$
×$\times(\mathbf{G^{\prime}-G})e^{i(\mathbf{G^{\prime}-G})\cdot\mathbf{R}_{I}}v_{\tau}^{\mathrm{NL}}(\mathbf{k+G,k+G^{\prime}}).$
Additionally, the stress tensor for the kinetic part and the nonlocal part are evaluated differently in sDFT and mDFT when compared to the traditinoal Kohn-Sham method. Specifically, the kinetic part is given by
@@
$=$ $\frac{2}{\Omega}\sum_{\mathbf{k}}W(\mathbf{k})\sum_{a,{\mathbf{GG^{\prime}}}}\gamma_{a\mathbf{k}}(\mathbf{G})(\mathbf{k+G})_{\alpha}\delta(\mathbf{GG^{\prime}})(\mathbf{k+G^{\prime}})_{\beta}\gamma_{a\mathbf{k}}(\mathbf{G^{\prime}}),$
and the nonlocal part is given by
$\sigma_{\alpha\beta}^{\mathrm{NL}}=$ $-\frac{2}{\Omega}{\rm Tr}\Big{[}\hat{f}_{H}\sum_{\mathbf{GG^{\prime}}\tau}S_{\tau}(\mathbf{G^{\prime}-G})\frac{\partial v_{\tau}^{\mathrm{NL}}(\mathbf{G+k,G^{\prime}+k})}{\partial\epsilon_{\alpha\beta}}|\mathbf{k+G}\rangle\langle\mathbf{k+G^{\prime}}|\Big{]}$ (66)
$=-\frac{2}{\Omega}\sum_{a{\mathbf{GG^{\prime}}},\tau}\gamma_{a\mathbf{k}}(\mathbf{G})S_{\tau}(\mathbf{G^{\prime}-G})\frac{\partial v_{\tau}^{\mathrm{NL}}(\mathbf{G+k,G^{\prime}+k})}{\partial\epsilon_{\alpha\beta}}\gamma_{a\mathbf{k}}(\mathbf{G^{\prime}}).$
Fig. 15(a) shows the parallel efficiency of sDFT and mDFT. sDFT demonstrates excellent scalability as all stochastic orbitals are independently evaluated, while mDFT involves calculating KS orbitals and additional data communication among CPU cores, reducing its efficiency compared to sDFT. Fig. 15(b) compares the operational efficiency at different temperatures by recording the average time per electronic iteration for the sDFT method. The temperature ranges from 5 to 300 eV, and 96 stochastic orbitals was adopted. In sDFT, the order of Chebyshev polynomial expansions was selected to ensure an electron error less than 1e-9. Specifically, for temperatures of 5, 10, 20, 30, 50, 100, 200, and 300 eV, Chebyshev orders of 1200, 620, 300, 220, 130, 60, 35, and 25 were used, respectively. As the temperature increases, the number of Chebyshev orders decreases, causing the wall time to drop exponentially. The sDFT and mDFT methods implemented in ABACUS has been successfully applied to study warm dense matter. Liu and Chen (2022b); Chen et al. (2024a)
Figure 15: Efficiency tests of the sDFT and mDFT methods. (a) Parallel efficiency of sDFT and mDFT for a C system ($N$=512) at the temperature of 20 eV. (b) Averaged wall time for an electronic iteration step of the sDFT method when calculating a C system ($N$=8) at a temperature range of 5-300 eV. All CPU cores are in Intel(R) Xeon(R) Platinum 9242 CPU @ 2.30GHz nodes.
IV.4 Orbital-Free DFT
IV.4.1 Formulas
@@
$L_{\rm{OF}}[\rho]=E_{\rm{OF}}[\rho]-\mu(\int{\rho(\mathbf{r})\text{d}^{3}\mathbf{r}}-N),$ (68)
@@
IV.4.2 Kinetic Energy Density Functional
@@
Several analytical KEDFs have been proposed over the past few decades, Karasiev and Trickey (2012); Witt et al. (2018) and they can be categorized into two main classes. First, the local and semilocal KEDFs are characterized by their kinetic energy density as a function of the charge density, its gradient, the Laplacian of the charge density, or even higher-order derivatives thereof. Thomas (1927); Fermi (1927); Weizsäcker (1935); Luo, Karasiev, and Trickey (2018); Constantin, Fabiano, and Della Sala (2018); Kang et al. (2020) Second, the nonlocal KEDFs define the kinetic energy density at each point in real space as a functional of the nonlocal charge density. Wang and Teter (1992); Wang, Govind, and Carter (1999); Huang and Carter (2010); Mi, Genova, and Pavanello (2018); Shao, Mi, and Pavanello (2021); Sun, Li, and Chen (2023); Bhattacharjee, Jana, and Samal (2024) Semilocal KEDFs are generally more computationally efficient, while nonlocal KEDFs tend to provide greater accuracy. However, a universally applicable KEDF that effectively describes both simple metal and semiconductor systems remains elusive, and a systematic approach to its construction has yet to be established.
@@
$T_{\rm{vW}}=\frac{1}{8}\int{\frac{{\left|{\nabla\rho({\mathbf{r}})}\right|}^{2}}{\rho\,({\mathbf{r}})}\,\text{d}^{3}{\mathbf{r}}}.$ (70)
@@
$T_{\rm{\theta}}=\int{\tau_{\rm{TF}}F_{\rm{\theta}}{\rm{d}}^{3}{\mathbf{r}}},$ (72)
@@
Notably, in a spin-degenerate system, the Pauli kinetic energy density can be expressed analytically using Kohn-Sham orbitals $\psi_{i}(\textbf{r})$ and corresponding occupation numbers $f_{i}$evy and Ou-Yang (1988)
@@
where $\varepsilon_{i}$ represents the eigenvalue associated with the Kohn-Sham orbital $\psi_{i}(\textbf{r})$. Additionally, $M$ denotes the highest occupied state, and $\varepsilon_{M}$ is the eigenvalue of $\psi_{M}(\textbf{r})$, which corresponds to the chemical potential $\mu$.
IV.4.3 Machine Learning Based Kinetic Energy Density Functional
In recent years, machine learning (ML) techniques have breathed new life into the development of KEDF. Snyder et al. (2012); Seino et al. (2018); Meyer, Weichselbaum, and Hauser (2020); Imoto, Imada, and Oshiyama (2021); Ryczko et al. (2022); Zhang et al. (2024a); Sun and Chen (2024) For example, Sun et al. imposed a machine learning based physical-constrained nonlocal (MPN) KEDF and implemented it in ABACUS. Sun and Chen (2024) The MPN KEDF is designed to satisfy three exact physical constraints: the scaling law of electron kinetic energy $T_{\rm{\theta}}[\rho_{\lambda}]=\lambda^{2}T_{\rm{\theta}}[\rho],\rho_{\lambda}=\lambda^{3}\rho(\lambda\textbf{r})$, the free electron gas (FEG) limit, and the non-negativity of Pauli energy density.
Figure 16: Workflow of the MPN KEDF. (Adapted with permission from Phys. Rev. B, 109(11): 115135 (2024). Copyright 2024 American Physical Society.)
@@ IV.4.3 Machine Learning Based Kinetic Energy Density Functional
$\tilde{p}_{\rm{nl}}(\mathbf{r})=\int{w(\mathbf{r}-\mathbf{r}^{\prime})\tilde{p}(\mathbf{r}^{\prime}){\rm{d}^{3}}\mathbf{r}^{\prime}},$ (82)
@@
$\tilde{\xi}(\mathbf{r})=\tanh{\left(\frac{\int{w(\mathbf{r}-\mathbf{r}^{\prime})\rho^{1/3}(\mathbf{r}^{\prime}){\rm{d}^{3}}\mathbf{r}^{\prime}}}{\rho^{1/3}(\mathbf{r})}\right)},\$ (84)
@@
$\tilde{\xi}_{\rm{nl}}(\mathbf{r})=\int{w(\mathbf{r}-\mathbf{r}^{\prime})\tilde{\xi}(\mathbf{r}^{\prime}){\rm{d}^{3}}\mathbf{r}^{\prime}}.$ (85)
@@
The training set for the MPN KEDF includes eight metallic structures, specifically bcc Li, fcc Mg, fcc Al, as well as five alloys: $\rm{Li_{3}Mg}$ (mp-976254), LiMg (mp-1094889), $\rm{Mg_{3}Al}$ (mp-978271), $\beta^{\prime\prime}$ $\rm{MgAl_{3}}$ Carling and Carter (2003), $\rm{LiAl_{3}}$ (mp-10890). The numbers in parentheses correspond to the Materials Project IDs. Jain et al. (2013)

### v3

@@ III.3.1 Formulas
The stochastic density functional theory (sDFT) Baer, Neuhauser, and Rabani (2013) was proposed to circumvent the $O(N^{3})$ scaling of KS-DFT calculations. Specifically, by introducing stochastic orbitals and the Chebyshev expansion scheme in the sDFT method, the electron density can be evaluated by tracing operators, bypassing the diagonalization of the Hamiltonian. Since the computational costs of tracing operations scale linearly with system size, this approach significantly enhances computational efficiency for large systems. Later, the finite-temperature sDFT Cytter et al. (2018) defines the Fermi-Dirac operator at finite temperature as
@@
where the chemical potential in $\hat{f}_{H}$ is determined by solving the conservation of electron number via
@@
The entropy is similarly formulated as:
$S=-2k_{B}\mathrm{Tr}\left[\hat{f}_{H}\ln\hat{f}_{H}+(1-\hat{f}_{H})\ln(1-\hat{f}_{H})\right].$ (43)
However, the use of stochastic orbitals inevitably introduces stochastic errors. To achieve higher accuracy, a large number of stochastic orbitals are typically required, which substantially increases the computational time despite its linear scaling property. For instance, fragmentation approaches have been developed to compute molecular systems Neuhauser, Baer, and Rabani (2014) and covalent materials Arnon et al. (2017); Chen et al. (2019a), while stochastic “embedding” methods have been used to calculate p-nitroaniline in water Li et al. (2019). Additionally, the energy window method proposed by Chen et al. Chen et al. (2019b) decomposes the electron density into different components based on orbital energies to reduce stochastic errors, demonstrating good generality and robustness.
For simulations of materials at extremely high temperatures, mixed stochastic-deterministic Density Functional Theory (mDFT) was proposed White and Collins (2020). The method combines the advantages of traditional KS-DFT and sDFT. The mDFT method retains the computational efficiency of sDFT for high-temperature systems while reducing stochastic errors using some of the deterministic Kohn-Sham orbitals.
III.3.2 Stochastic Orbitals
@@
$\langle\phi_{j}|\chi_{a}\rangle=\frac{1}{\sqrt{N_{\chi}}}\exp(i2\pi\theta_{j}^{a}),$ (44)
@@
[…]
@@
For example, the formula for electron density in mDFT is
@@
Note that here we use a finite number of stochastic orbitals $N_{\chi}$, which introduces some stochastic errors.
III.3.3 Plane-Wave-Based Implementation
We have implemented the sDFT and mDFT methods based on plane wave basis Liu and Chen (2022b) and periodic boundary conditions in ABACUS. Furthermore, both methods can be used with the $k$-point sampling method. The flowchart of sDFT and mDFT is shown in Fig. 13.
Figure 13: Flowchart of stochastic DFT in ABACUS. When the number of KS orbitals $N_{\phi}$ is larger than 0, it is the flowchart of mixed stochastic-deterministic DFT. When the number of KS orbitals $N_{\phi}$ reduces to 0, indicating that the dashed box does not exist, it degenerates into the flowchart of sDFT. The chemical potential is $\mu$.
@@ III.3.3 Plane-Wave-Based Implementation
$\rho(\mathbf{r})=2\sum_{\mathbf{k}}W(\mathbf{k})\sum_{a\mathbf{G}}\left|\left\langle\mathbf{r}|\mathbf{k+G}\right\rangle\left\langle\mathbf{k+G}\left|\hat{f}_{H}^{1/2}\right|\chi_{a\mathbf{k}}\right\rangle\right|^{2}.$ (53)
Let us define
@@
so the electron density can be written as
@@
For the atomic forces, only the nonlocal potential part is different from the traditional Kohn-Sham method, which is given by
$\mathbf{F}_{I}^{\mathrm{NL}}=$ $2{\rm Tr}\Big{[}-i\hat{f}_{H}\sum_{\mathbf{GG^{\prime}}}(\mathbf{G^{\prime}-G})e^{i(\mathbf{G^{\prime}-G})\cdot\mathbf{R}_{I}}$ (56)
@@
$=$ $-2i\sum_{\mathbf{k}}W(\mathbf{k})\sum_{a{\mathbf{GG^{\prime}}}}\gamma_{a\mathbf{k}}(\mathbf{G})\gamma_{a\mathbf{k}}(\mathbf{G^{\prime}})$
$\times(\mathbf{G^{\prime}-G})e^{i(\mathbf{G^{\prime}-G})\cdot\mathbf{R}_{I}}v_{\tau}^{\mathrm{NL}}(\mathbf{k+G,k+G^{\prime}}).$
Additionally, the stress tensor for the kinetic part and the nonlocal part is evaluated differently in sDFT and mDFT when compared to the traditional Kohn-Sham method. Specifically, the kinetic part is
@@
$=$ $\frac{2}{\Omega}\sum_{\mathbf{k}}W(\mathbf{k})\sum_{a{\mathbf{GG^{\prime}}}}\gamma_{a\mathbf{k}}(\mathbf{G})(\mathbf{k+G})_{\alpha}\delta(\mathbf{GG^{\prime}})(\mathbf{k+G^{\prime}})_{\beta}\gamma_{a\mathbf{k}}(\mathbf{G^{\prime}}),$
and the nonlocal part takes the form of
$\sigma_{\alpha\beta}^{\mathrm{NL}}$ $=-\frac{2}{\Omega}{\rm Tr}\Big{[}\hat{f}_{H}\sum_{\mathbf{GG^{\prime}}\tau}S_{\tau}(\mathbf{G^{\prime}-G})$ (58)
$\qquad\quad\frac{\partial v_{\tau}^{\mathrm{NL}}(\mathbf{G+k,G^{\prime}+k})}{\partial\epsilon_{\alpha\beta}}|\mathbf{k+G}\rangle\langle\mathbf{k+G^{\prime}}|\Big{]}$
$=-\frac{2}{\Omega}\sum_{a{\mathbf{GG^{\prime}}}\tau}\gamma_{a\mathbf{k}}(\mathbf{G})S_{\tau}(\mathbf{G^{\prime}-G})$
$\qquad\quad\frac{\partial v_{\tau}^{\mathrm{NL}}(\mathbf{G+k,G^{\prime}+k})}{\partial\epsilon_{\alpha\beta}}\gamma_{a\mathbf{k}}(\mathbf{G^{\prime}}).$
Fig. 14(a) shows the parallel efficiency of sDFT and mDFT. sDFT demonstrates excellent scalability as all stochastic orbitals are independently evaluated, while mDFT involves calculating KS orbitals and additional data communication among CPU cores, reducing its efficiency compared to sDFT. Fig. 14(b) compares the operational efficiency at different temperatures by recording the average time per electronic iteration for the sDFT method. The temperature ranges from 5 to 300 eV, and 96 stochastic orbitals were adopted. In sDFT, the order of Chebyshev polynomial expansions was selected to ensure an electron error less than 1e-9. Specifically, for temperatures of 5, 10, 20, 30, 50, 100, 200, and 300 eV, Chebyshev orders of 1200, 620, 300, 220, 130, 60, 35, and 25 were used, respectively. As the temperature increases, the number of Chebyshev orders decreases, causing the wall time to drop exponentially. The sDFT and mDFT methods implemented in ABACUS has been successfully applied to study warm dense matter. Liu and Chen (2022b); Chen et al. (2024a, 2025)
Figure 14: Efficiency tests of the sDFT and mDFT methods. (a) Parallel efficiency of sDFT and mDFT for a C system with the number of atoms being NN=512 at the temperature of 20 eV. (b) Averaged wall time for an electronic iteration step of the sDFT method when calculating a C system (NN=8) at a temperature range from 5 to 300 eV. The tested machines are Intel(R) Xeon(R) Platinum 9242 CPU @ 2.30GHz nodes.
III.4 Orbital-Free DFT
III.4.1 Formulas
@@
$L_{\rm{OF}}[\rho]=E_{\rm{OF}}[\rho]-\mu(\int{\rho(\mathbf{r})\text{d}\mathbf{r}}-N),$ (60)
@@
III.4.2 Kinetic Energy Density Functional
@@
Several analytical KEDFs have been proposed over the past few decades, Karasiev and Trickey (2012); Witt et al. (2018) and they can be categorized into two main classes. First, the local and semilocal KEDFs are characterized by their kinetic energy density as a function of the charge density, its gradient, the Laplacian of the charge density, or even higher-order derivatives thereof. Thomas (1927); Fermi (1927); Weizsäcker (1935); Luo, Karasiev, and Trickey (2018); Constantin, Fabiano, and Della Sala (2018); Kang et al. (2020) Second, the nonlocal KEDFs define the kinetic energy density at each point in real space as a functional of the nonlocal charge density. Wang and Teter (1992); Wang, Govind, and Carter (1999); Huang and Carter (2010); Mi, Genova, and Pavanello (2018); Shao, Mi, and Pavanello (2021); Sun, Li, and Chen (2023); Bhattacharjee, Jana, and Samal (2024) Semilocal KEDFs are generally more computationally efficient, while nonlocal KEDFs tend to provide greater accuracy. However, a universally applicable KEDF that effectively describes both simple metals and semiconductor systems remains largely elusive, and a systematic approach to its development remains to be explored.
@@
$T_{\rm{vW}}=\frac{1}{8}\int{\frac{{\left|{\nabla\rho({\mathbf{r}})}\right|}^{2}}{\rho\,({\mathbf{r}})}\,\text{d}{\mathbf{r}}}.$ (62)
@@
$T_{\rm{\theta}}=\int{\tau_{\rm{TF}}F_{\rm{\theta}}{\rm{d}}{\mathbf{r}}},$ (64)
@@
Notably, in a spin-degenerate system, the Pauli kinetic energy density can be expressed analytically using Kohn-Sham orbitals $\psi_{i}(\textbf{r})$ and corresponding occupation numbers $f_{i}$Levy and Ou-Yang (1988)
@@
where $\varepsilon_{i}$ represents the eigenvalue associated with the Kohn-Sham orbital $\psi_{i}(\textbf{r})$. Additionally, $M$ denotes the highest occupied state, and $\varepsilon_{M}$ is the eigenvalue of $\psi_{M}(\textbf{r})$, which corresponds to the chemical potential $E_{\mu}$.
III.4.3 Machine Learning Based Kinetic Energy Density Functional
In recent years, machine learning (ML) techniques have breathed new life into the development of KEDF. Snyder et al. (2012); Seino et al. (2018); Meyer, Weichselbaum, and Hauser (2020); Imoto, Imada, and Oshiyama (2021); Ryczko et al. (2022); Zhang et al. (2024a); Sun and Chen (2024a) For example, Sun et al. imposed a machine learning based physical-constrained nonlocal (MPN) KEDF and implemented it in ABACUS. Sun and Chen (2024a) The MPN KEDF is designed to satisfy three exact physical constraints: the scaling law of electron kinetic energy $T_{\rm{\theta}}[\rho_{\lambda}]=\lambda^{2}T_{\rm{\theta}}[\rho],\rho_{\lambda}=\lambda^{3}\rho(\lambda\textbf{r})$, the free electron gas (FEG) limit, and the non-negativity of Pauli energy density.
Figure 15: Workflow of the MPN kinetic energy density functional within the framework of orbital-free DFT. (Adapted with permission from Phys. Rev. B, 109(11): 115135 (2024). Copyright 2024 American Physical Society.)
@@ III.4.3 Machine Learning Based Kinetic Energy Density Functional
$\tilde{p}_{\rm{nl}}(\mathbf{r})=\int{w(\mathbf{r}-\mathbf{r}^{\prime})\tilde{p}(\mathbf{r}^{\prime}){\rm{d}}\mathbf{r}^{\prime}},$ (74)
@@
[…]
@@
$\tilde{\xi}_{\rm{nl}}(\mathbf{r})=\int{w(\mathbf{r}-\mathbf{r}^{\prime})\tilde{\xi}(\mathbf{r}^{\prime}){\rm{d}}\mathbf{r}^{\prime}}.$ (77)
@@
While the MPN KEDF has shown promising results for simple metals and their alloys Sun and Chen (2024a), its performance remains limited for semiconductors. To address this, Sun et al. developed a multi-channel MPN KEDF Sun and Chen (2024b) that incorporates information from multiple real-space length scales.
IV Methods in Numerical Atomic Orbital Basis
IV.1 Numerical Atomic Orbitals
The efficiency and accuracy of first-principles methods are largely determined by the basis sets. As mentioned in Sec. III, although the accuracy of the plane wave basis in solving Kohn-Sham equation can be systematically improved by increasing the kinetic energy cutoff, its computational costs are generally formidable for systems consisting of hundreds or thousands of atoms. On the other hand, atomic orbital basis sets are generally more efficient than plane wave basis sets, at the cost of a slight loss in accuracy. In particular, numerical atomic orbitals (NAOs) have become a competitive option for basis set types, especially for calculating large systems consisting of hundreds or even thousands of atoms. The NAO basis set also supports the linear scaling algorithms due to the strict locality feature. Over the past several decades, several methods have been proposed Talman (1978); Sharafeddin et al. (1992); Toyoda and Ozaki (2010); Cerioni et al. (2012); Soler et al. (2002); Blum et al. (2009) to perform efficient numerical integrations for NAOs.
Unlike the plane-wave basis, there is no unique way to construct NAOs. For example, a popular method involves solving isolated atoms subject to confining potentials Sankey and Niklewski (1989); Porezag et al. (1995); Horsfield (1997); Junquera et al. (2001); Soler et al. (2002). Based on confining potentials, Blum et alBlum et al. (2009) proposed to generate NAOs by iteratively picking up basis functions one by one from a pool of predefined candidates, to seek the best improvement of a target energy. Alternatively, OzakiOzaki (2003); Ozaki and Kino (2004) suggested that eigenfunctions of isolated atoms in confining potentials serve as “primitive orbitals” in terms of which NAOs are expanded, and expansion coefficients can be optimized with self-consistent calculations. Apart from the above energy-based methods, another class of methods constructs NAOs towards some reference states. For example, Sanchez-Portal et al Sanchez-Portal, Artacho, and Soler (1995); Sánchez-Portal, Artacho, and Soler (1996) proposed to optimize the following “spillage” function defined as